7. A probability generating function encodes the probabilities of different events happening with a gener- ating function. For example, consider tossing a coin until the first time you get a heads. Then, let pn be the probability that you had to flip the coin n times until you got a heads for the first time. The probability generating function for this problem would be f(z) = Σ Pnz. 膨:1 (a) If your coin has a probability p of landing heads on any particular coin toss, find the value of pa in terms of p. (b) Find a closed form for the probability generating function for this coin (c) What is f(1)? Give a non-algebraic reason for why this is true. (d) The average number of times you need to flip the coin in order to get a heads would be given by n Pn Compute this value using your generating function from part (b)
Answer #1

(a)

For getting heads for the first time on nth toss, we have got all tails on (n-1) tosses.

probability of tails = 1-p and all tosses are independent of each other, then

p_n = (1-p)^{n-1}p

(b)

n-1 アn n= 1 7l n= 1 7l n= 1

= pxsum_{n=0}^{infty } ((1-p)x)^{n}

= px rac{1}{1 - (1-p)x} = rac{px}{1 - x + px} (By sum of geometric series with common ratio = (1-p)x and first term = 1}

(c)

1-1+p*1=p/p = 1

This is true because f(1) is sum of all probabilities for possible values of n which is always equal to 1.

(d)

By definition of probability generating function, average number of times need to flip the coin is,

f'(1)= rac{mathrm{d} }{mathrm{d} x}_{x=1} rac{px}{1 - x + px} = rac{mathrm{d} }{mathrm{d} x}_{x=1} rac{p(1-x+px)- px(-1 + p)}{(1 - x + px)^2}

= rac{p(1-1+p*1)- p*1(-1 + p)}{(1 - 1 + p*1)^2} = rac{p^2 + p - p^2}{p^2} = rac{1}{p}